# HOLO: Homography-Guided Pose Estimator Network for Fine-Grained Visual Localization on SD Maps

### v1

@@ 1 Introduction
Figure 1: Comparision of attention-based direct pose estimation and homography-guided pose estimation. (a) Schematic diagram of attention-based direct pose estimation and homography-guided pose estimation. (b) The evolution of localization error across training iterations between two methods on nuScene[caesar2020nuscenes].
Recent studies have explored diverse map representations for localization, including satellite maps [lentsch2023slicematch, wang2023fine, xia2023convolutional, yuan2024cross], high-definition (HD) maps [he2024egovm, zhang2025bev, Zhao2024towards], and 3D prior maps [chen2024map, Dong_2025_CVPR]. In autonomous driving scenarios, early research on visual localization mainly relied on HD maps. However, HD maps are costly to construct, require frequent maintenance, and lack scalability. These limitations have motivated growing interest in using lightweight and easily accessible standard-definition maps (SD) maps, such as OpenStreetMap [Haklay2008open], for large-scale localization.
Early works [samano2020you, zhou2021efficient] formulated visual localization as an image retrieval task to perform coarse localization. With the rise of bird’s-eye-view (BEV) perception[philion2020lift, li2024bevformer, li2023fb], research has shifted toward geometric alignment between BEV features and map priors for improved accuracy. OrienterNet [sarlin2023orienternet] introduced an end-to-end framework performing exhaustive pose-wise BEV-to-map matching but suffered from high computational cost. To enhance efficiency, MapLocNet [wu2024maplocnet] employs a regression-based approach to directly predict the target pose, avoiding the complex matching process. However, while matching methods only need to distinguish similarities within a limited set of pose candidates, regression requires fitting continuous pose values, which makes training more challenging and convergence slower.
To address this problem, we identify two main limitations in existing regression-based approaches. First, during feature fusion, the absence of explicit geometric guidance leads most methods [wu2024maplocnet, he2024egovm, zhang2025bev] to rely solely on attention-based mechanisms to learn cross-modal correlations, which often results in low efficiency. Second, directly regressing a 3-DoF pose without geometric constraints introduces unstable gradients, making optimization challenging and prone to overfitting.
@@
To promote further research on visual localization with SD maps, we collected OpenStreetMap (OSM) data corresponding to the four cities in the nuScenes[caesar2020nuscenes] dataset, which we then aligned with the provided HD maps and plan to release publicly. Experimental results demonstrate that our approach significantly outperforms the current state-of-the-art visual localization method. The main contributions of this work are summarized as follows:
@@ 2.1 Visual Localization with Maps
Matching-based methods. These methods formulate image-to-map localization as a pose retrieval problem, where the goal is to learn discriminative embeddings for both images and maps and maximize their similarity at the corresponding poses. Some works [lentsch2023slicematch, xia2023convolutional] focus on encoding ground-view and aerial-view images into pose-aware feature vectors that jointly capture spatial and orientation information for retrieval, while others [sarlin2023orienternet, sarlin2023snap] extend this idea by learning implicit BEV features and projecting them onto the map according to pose candidates to compute matching scores, where the highest-scoring hypothesis is selected as the final pose. Based on this framework, SegLocNet[zhou2025seglocnet] further replaces implicit features with binarized semantic masks to improve localization precision. Although these methods achieve strong performance through representation learning, they exhibit inherent limitations. Since they rely on discrete pose sampling for feature matching, their localization accuracy is constrained by the sampling resolution. Moreover, as the number of pose hypotheses increases, the inference time grows significantly, resulting in a trade-off between accuracy and efficiency.
Regression-based methods. Unlike matching-based approaches that require pose discretization, regression-based methods infer the camera pose directly from paired image and map inputs. These approaches generally comprise three core components: feature extraction, cross-modal fusion, and pose decoding. Recent approaches [chen2024map, wu2024maplocnet] leverage transformer-based encoders for effective feature fusion, whereas subsequent studies [Dong_2025_CVPR, yuan2024cross, zhang2025bev] further enhance inter-modal interactions through cross-attention mechanisms. Most of these methods focus on designing robust fusion modules to obtain features that can be more easily mapped to pose, thereby improving localization accuracy. However, they mainly relied on attention mechanisms to learn feature associations implicitly, without incorporating any geometric priors into this stage. This often led to low efficiency and poor accuracy.
@@ 2.2 Cross-modal Homography Estimation
Homography estimation aims to recover a planar projective transformation between two views by estimating a 3x3 homography matrix that maps corresponding points across image planes. In cross-modal settings, the two inputs originate from different sensing modalities or data sources, such as RGB vs. NIR or optical vs. SAR imagery. Early deep homography methods [zhao2021deep, cao2022iterative, cao2023recurrent] extended supervised homography learning to the cross-modal case under strong direct supervision. However, ground-truth homography annotations are rarely available in real-world scenarios, which significantly limits the practicality of such approaches.
To alleviate this constraint, recent works explore unsupervised cross-modal homography estimation. Zhang et al. [zhang2024scpnet] align heterogeneous image pairs through a consistent feature projection module and leverage intra-modal self-supervision to compensate for missing ground-truth labels. Song et al. [song2024unsupervised] alternate between modality adaptation and geometric refinement to progressively reduce appearance and alignment discrepancies. Yu et al. [yu2025sshnet] further decomposes the problem into modality alignment and homography regression, optimizing the two sub-networks separately to stabilize training.
Inspired by this decomposition strategy, we formulate multi-view image to SD map localization in a similar manner: we first reduce the modality gap between images and SD maps using a semantic alignment network, and then regress the relative pose using a homography estimation network. Unlike [yu2025sshnet], which requires stage-wise training due to the absence of geometric supervision, our task benefits from available GT vehicle poses, allowing us to jointly optimize both components in an end-to-end manner.
@@ 3.2 BEV Perception Module
Our BEV perception module aims to project the surround-view images $I$ into the BEV space to reduce the perspective gap between camera observations and the map for downstream homography estimation tasks. Following prior works such as LSS [philion2020lift] and OrienterNet [sarlin2023orienternet], we adopt a depth-guided projection mechanism that transforms image features into BEV representations in a simple yet effective manner.
Camera Feature Extraction: We utilized a lightweight pre-trained EfficientNet-B0[tan2019efficientnet] backbone to extract multi-scale semantic features from all six ground images. To model scene geometry, the network jointly predicts a discrete depth distribution and semantic embeddings through a shared head, generating depth-aware volumetric features.
@@ 3.3 Map Process Module
Map Encoding: Similar to OrienterNet [sarlin2023orienternet], we first assign an $N$-dimensional learnable embedding to each class in the rasterized map, producing a feature map of size $H\times W\times 2N$. The embedded map is then processed by a U-Net architecture built upon VGG16 [vgg] to extract semantic features. The resulting map feature $\mathbf{\Phi}_{map}\in\mathbb{R}^{C\times H_{map}\times W_{map}}$ maintains the same size as the BEV features $\mathbf{\Phi}_{bev}$. The map feature is also passed through a semantic segmentation head to produce the corresponding map mask $\mathbf{M}_{map}^{sem}\in\mathbb{R}^{2\times H_{map}\times W_{map}}$.
@@ 3.4 Homography-Guided Pose Estimation Module
In this module, we introduce a high–efficiency pose regression framework. Inspired by IHN[cao2022iterative], rather than relying on attention mechanisms to implicitly infer cross–modal correspondences, we adopt an explicit strategy: using the estimated homography to guide the correlation formation between the BEV and the map.
Feature Fusion. We use a Siamese ResNet[he2016deep] encoder to extract features for homography estimation. The input semantic features $\mathbf{\Phi}_{BEV}$ and $\mathbf{\Phi}_{Map}$ are downsampled by $4\times$ and projected to homography features via a $1\times 1$ convolution. We denote these two features as $\mathbf{F}_{BEV}\in\mathbb{R}^{D\times H^{\prime}\times W^{\prime}}$ and $\mathbf{F}_{Map}\in\mathbb{R}^{D\times H^{\prime}\times W^{\prime}}$, where $H^{\prime},W^{\prime}=\frac{H}{4},\frac{W}{4}$. To get fused feature, We form a dense correlation volume $\mathbf{C}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times H^{\prime}\times W^{\prime}}$ by applying dot-product matching between all spatial locations of $\mathbf{F}_{BEV}$ and $\mathbf{F}_{Map}$:
@@
The fused features $\mathbf{S}_{k}$ and $\mathbf{S}^{1/2}_{k}$ are concatenated and fed into the pose decoder. Instead of directly predicting the 3-DoF pose $(x,y,\theta)$, the pose decoder outputs the corner displacement $\mathbf{\Delta D}_{k}$ for the current iteration. The estimated displacement $\mathbf{\Delta D}_{k}$ is then converted into the homography matrix $\mathbf{H}_{k}$ via Direct Linear Transformation (DLT)[abdel2015direct], which is used to update the projected coordinates $\mathbf{X}$ for the next iteration. Moreover, the estimated $\mathbf{H}_{k}$ further induces a geometric mapping allowing us to recover the pose parameters at this iteration.
@@ 3.5 Mapping 3D Pose by Homography matrix
We replace direct pose regression with a homography-based projection strategy. This approach constrains the pose within a feasible solution space, thereby improving estimation accuracy and convergence speed. The homography establishes the pixel-wise mapping between two planes as
@@ 4 Experiments
SegLocNet-road*[zhou2025seglocnet] $C+L$ $35.63$ $57.98$ $74.55$ $80.09$ $38.58$ $64.98$ $83.62$ $87.40$ $8.15$ $19.68$
SegLocNet-drivable*[zhou2025seglocnet] $C+L$ $59.08$ $76.04$ $84.25$ $86.86$ $63.19$ $84.55$ $91.02$ $93.00$ 5.30{5.30} 10.11{10.11}
OrienterNet[sarlin2023orienternet] $C$ $15.74$ $33.27$ $49.86$ $60.10$ $31.69$ $45.81$ $61.23$ $72.44$ $15.47$ $26.43$
U-BEV*[camiletto2024u] $C$ $16.89$ $41.60$ $71.33$ $83.46$ - - - - - -
MapLocNet One-Stage*[wu2024maplocnet] $C$ $16.32$ $40.56$ $74.27$ $89.69$ $53.71$ $78.13$ $94.49$ $98.05$ - -
MapLocNet*[wu2024maplocnet] $C$ $20.10$ $45.54$ $77.70$ $91.89$ $58.61$ $84.10$ $96.23$ $98.62$ - -
@@ 4.1 Datasets
We conduct training and evaluation on the nuScenes[caesar2020nuscenes] autonomous driving dataset. NuScenes contains 1000 real-world driving scenes collected across Boston and Singapore. Following the official data split, we adopt 850 scenes for training and use the remaining 150 scenes as the validation set to benchmark our model’s performance.
Since nuScenes does not provide SD maps, we retrieve the corresponding regions from OSM and align them to the nuScenes coordinate system. Unlike prior work [Blos-BEV, wu2024maplocnet, zhou2025seglocnet] that aligns OSM maps in EPSG:3857 and requires manual scaling for the Boston area, we directly transform WGS84 coordinates into the ENU frame used by nuScenes. Specifically, we choose a fixed reference origin per region, convert WGS84 to ECEF, and then map ECEF to a local ENU frame. A small constant drift correction is applied to ensure precise alignment, particularly for the Singapore scenes where minor GPS drift exists in the raw data. As a results, well-aligned SD maps were obtained for each keyframe across 1,000 scenes, thereby extending the nuScenes dataset.
@@ 4.2 Implementation Details
Training details. We implement our network in PyTorch. Training is performed on an NVIDIA RTX A6000 GPU using the AdamW[loshchilov2017decoupled] optimizer with a maximum learning rate of $3.5\times 10^{-4}$ and a weight decay of $5\times 10^{-4}$. The batch size is set to 16, and the model is trained for 180,000 iterations. A OneCycle[smith2017cyclical] learning rate scheduler is employed to adjust the learning rate throughout training. The loss weights are set as: $\lambda_{\text{sem}}=1000,\lambda_{\text{trans}}=1,\lambda_{\text{ori}}=10$ and the iterative pose refinement is performed for $N=6$ iterations.
Metrics. We follow the evaluation metrics defined in [zhou2025seglocnet, wu2024maplocnet], including Recall@X m, Recall@X°, Absolute Position Error (APE), and Absolute Orientation Error (AOE), which jointly assess translational and rotational localization accuracy.
@@ 4.3 Comparisons with Existing Methods
We compare our proposed network HOLO with several existing methods: OrienterNet [sarlin2023orienternet], U-BEV [camiletto2024u], MapLocNet [wu2024maplocnet], SegLocNet [zhou2025seglocnet]. In addition, to further validate the effectiveness of our homography-guided strategy, we implement a pose estimation network based on cross-attention feature fusion[sun2021loftr, qin2022geometric], namely HOLO-CA, as shown in Fig. 4. This network is also included in our comparison.

### v2

@@ 1 Introduction
Figure 1: Comparision of attention-based direct pose estimation and homography-guided pose estimation. (a) Schematic diagram of attention-based direct pose estimation and homography-guided pose estimation. (b) The evolution of localization error across training iterations between two methods on nuScene[2].
Recent studies have explored diverse map representations for localization, including satellite maps [11, 25, 28, 30], high-definition (HD) maps [10, 32, 33], and 3D prior maps [6, 7]. In autonomous driving scenarios, early research on visual localization mainly relied on HD maps. However, HD maps are costly to construct, require frequent maintenance, and lack scalability. These limitations have motivated growing interest in using lightweight and easily accessible standard-definition maps (SD) maps, such as OpenStreetMap [8], for large-scale localization.
Early works [17, 35] formulated visual localization as an image retrieval task to perform coarse localization. With the rise of bird’s-eye-view (BEV) perception[15, 12, 13], research has shifted toward geometric alignment between BEV features and map priors for improved accuracy. OrienterNet [18] introduced an end-to-end framework performing exhaustive pose-wise BEV-to-map matching but suffered from high computational cost. To enhance efficiency, MapLocNet [26] employs a regression-based approach to directly predict the target pose, avoiding the complex matching process. However, while matching methods only need to distinguish similarities within a limited set of pose candidates, regression requires fitting continuous pose values, which makes training more challenging and convergence slower.
To address this problem, we identify two main limitations in existing regression-based approaches. First, during feature fusion, the absence of explicit geometric guidance leads most methods [26, 10, 32] to rely solely on attention-based mechanisms to learn cross-modal correlations, which often results in low efficiency. Second, directly regressing a 3-DoF pose without geometric constraints introduces unstable gradients, making optimization challenging and prone to overfitting.
@@
To promote further research on visual localization with SD maps, we collected OpenStreetMap (OSM) data corresponding to the four cities in the nuScenes[2] dataset, which we then aligned with the provided HD maps and plan to release publicly. Experimental results demonstrate that our approach significantly outperforms the current state-of-the-art visual localization method. The main contributions of this work are summarized as follows:
@@ 2.1 Visual Localization with Maps
Matching-based methods. These methods formulate image-to-map localization as a pose retrieval problem, where the goal is to learn discriminative embeddings for both images and maps and maximize their similarity at the corresponding poses. Some works [11, 28] focus on encoding ground-view and aerial-view images into pose-aware feature vectors that jointly capture spatial and orientation information for retrieval, while others [18, 19] extend this idea by learning implicit BEV features and projecting them onto the map according to pose candidates to compute matching scores, where the highest-scoring hypothesis is selected as the final pose. Based on this framework, SegLocNet[36] further replaces implicit features with binarized semantic masks to improve localization precision. Although these methods achieve strong performance through representation learning, they exhibit inherent limitations. Since they rely on discrete pose sampling for feature matching, their localization accuracy is constrained by the sampling resolution. Moreover, as the number of pose hypotheses increases, the inference time grows significantly, resulting in a trade-off between accuracy and efficiency.
Regression-based methods. Unlike matching-based approaches that require pose discretization, regression-based methods infer the camera pose directly from paired image and map inputs. These approaches generally comprise three core components: feature extraction, cross-modal fusion, and pose decoding. Recent approaches [6, 26] leverage transformer-based encoders for effective feature fusion, whereas subsequent studies [7, 30, 32] further enhance inter-modal interactions through cross-attention mechanisms. Most of these methods focus on designing robust fusion modules to obtain features that can be more easily mapped to pose, thereby improving localization accuracy. However, they mainly relied on attention mechanisms to learn feature associations implicitly, without incorporating any geometric priors into this stage. This often led to low efficiency and poor accuracy.
@@ 2.2 Cross-modal Homography Estimation
Homography estimation aims to recover a planar projective transformation between two views by estimating a 3x3 homography matrix that maps corresponding points across image planes. In cross-modal settings, the two inputs originate from different sensing modalities or data sources, such as RGB vs. NIR or optical vs. SAR imagery. Early deep homography methods [34, 4, 5] extended supervised homography learning to the cross-modal case under strong direct supervision. However, ground-truth homography annotations are rarely available in real-world scenarios, which significantly limits the practicality of such approaches.
To alleviate this constraint, recent works explore unsupervised cross-modal homography estimation. Zhang et al. [31] align heterogeneous image pairs through a consistent feature projection module and leverage intra-modal self-supervision to compensate for missing ground-truth labels. Song et al. [22] alternate between modality adaptation and geometric refinement to progressively reduce appearance and alignment discrepancies. Yu et al. [29] further decomposes the problem into modality alignment and homography regression, optimizing the two sub-networks separately to stabilize training.
Inspired by this decomposition strategy, we formulate multi-view image to SD map localization in a similar manner: we first reduce the modality gap between images and SD maps using a semantic alignment network, and then regress the relative pose using a homography estimation network. Unlike [29], which requires stage-wise training due to the absence of geometric supervision, our task benefits from available GT vehicle poses, allowing us to jointly optimize both components in an end-to-end manner.
@@ 3.2 BEV Perception Module
Our BEV perception module aims to project the surround-view images $I$ into the BEV space to reduce the perspective gap between camera observations and the map for downstream homography estimation tasks. Following prior works such as LSS [15] and OrienterNet [18], we adopt a depth-guided projection mechanism that transforms image features into BEV representations in a simple yet effective manner.
Camera Feature Extraction: We utilized a lightweight pre-trained EfficientNet-B0[24] backbone to extract multi-scale semantic features from all six ground images. To model scene geometry, the network jointly predicts a discrete depth distribution and semantic embeddings through a shared head, generating depth-aware volumetric features.
@@ 3.3 Map Process Module
Map Encoding: Similar to OrienterNet [18], we first assign an $N$-dimensional learnable embedding to each class in the rasterized map, producing a feature map of size $H\times W\times 2N$. The embedded map is then processed by a U-Net architecture built upon VGG16 [20] to extract semantic features. The resulting map feature $\mathbf{\Phi}_{map}\in\mathbb{R}^{C\times H_{map}\times W_{map}}$ maintains the same size as the BEV features $\mathbf{\Phi}_{bev}$. The map feature is also passed through a semantic segmentation head to produce the corresponding map mask $\mathbf{M}_{map}^{sem}\in\mathbb{R}^{2\times H_{map}\times W_{map}}$.
@@ 3.4 Homography-Guided Pose Estimation Module
In this module, we introduce a high–efficiency pose regression framework. Inspired by IHN[4] and HCNet[25], rather than relying on attention mechanisms to implicitly infer cross–modal correspondences, we adopt an explicit strategy: using the estimated homography to guide the correlation formation between the BEV and the map.
Feature Fusion. We use a Siamese ResNet[9] encoder to extract features for homography estimation. The input semantic features $\mathbf{\Phi}_{BEV}$ and $\mathbf{\Phi}_{Map}$ are downsampled by $4\times$ and projected to homography features via a $1\times 1$ convolution. We denote these two features as $\mathbf{F}_{BEV}\in\mathbb{R}^{D\times H^{\prime}\times W^{\prime}}$ and $\mathbf{F}_{Map}\in\mathbb{R}^{D\times H^{\prime}\times W^{\prime}}$, where $H^{\prime},W^{\prime}=\frac{H}{4},\frac{W}{4}$. To get fused feature, We form a dense correlation volume $\mathbf{C}\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times H^{\prime}\times W^{\prime}}$ by applying dot-product matching between all spatial locations of $\mathbf{F}_{BEV}$ and $\mathbf{F}_{Map}$:
@@
The fused features $\mathbf{S}_{k}$ and $\mathbf{S}^{1/2}_{k}$ are concatenated and fed into the pose decoder. Instead of directly predicting the 3-DoF pose $(x,y,\theta)$, the pose decoder outputs the corner displacement $\mathbf{\Delta D}_{k}$ for the current iteration. The estimated displacement $\mathbf{\Delta D}_{k}$ is then converted into the homography matrix $\mathbf{H}_{k}$ via Direct Linear Transformation (DLT)[1], which is used to update the projected coordinates $\mathbf{X}$ for the next iteration. Moreover, the estimated $\mathbf{H}_{k}$ further induces a geometric mapping allowing us to recover the pose parameters at this iteration.
@@ 3.5 Mapping 3D Pose by Homography matrix
We replace direct pose regression with a homography-based projection strategy. The projection process is similar to [25]. But, interestingly, we find that this approach can constrain the pose within a feasible solution space, thereby improving estimation accuracy and convergence speed in our work. The homography establishes the pixel-wise mapping between two planes as
@@ 4 Experiments
SegLocNet-road*[36] $C+L$ $35.63$ $57.98$ $74.55$ $80.09$ $38.58$ $64.98$ $83.62$ $87.40$ $8.15$ $19.68$
SegLocNet-drivable*[36] $C+L$ $59.08$ $76.04$ $84.25$ $86.86$ $63.19$ $84.55$ $91.02$ $93.00$ 5.30{5.30} 10.11{10.11}
OrienterNet[18] $C$ $15.74$ $33.27$ $49.86$ $60.10$ $31.69$ $45.81$ $61.23$ $72.44$ $15.47$ $26.43$
U-BEV*[3] $C$ $16.89$ $41.60$ $71.33$ $83.46$ - - - - - -
MapLocNet One-Stage*[26] $C$ $16.32$ $40.56$ $74.27$ $89.69$ $53.71$ $78.13$ $94.49$ $98.05$ - -
MapLocNet*[26] $C$ $20.10$ $45.54$ $77.70$ $91.89$ $58.61$ $84.10$ $96.23$ $98.62$ - -
@@ 4.1 Datasets
We conduct training and evaluation on the nuScenes[2] autonomous driving dataset. NuScenes contains 1000 real-world driving scenes collected across Boston and Singapore. Following the official data split, we adopt 850 scenes for training and use the remaining 150 scenes as the validation set to benchmark our model’s performance.
Since nuScenes does not provide SD maps, we retrieve the corresponding regions from OSM and align them to the nuScenes coordinate system. Unlike prior work [27, 26, 36] that aligns OSM maps in EPSG:3857 and requires manual scaling for the Boston area, we directly transform WGS84 coordinates into the ENU frame used by nuScenes. Specifically, we choose a fixed reference origin per region, convert WGS84 to ECEF, and then map ECEF to a local ENU frame. A small constant drift correction is applied to ensure precise alignment, particularly for the Singapore scenes where minor GPS drift exists in the raw data. As a results, well-aligned SD maps were obtained for each keyframe across 1,000 scenes, thereby extending the nuScenes dataset.
@@ 4.2 Implementation Details
Training details. We implement our network in PyTorch. Training is performed on an NVIDIA RTX A6000 GPU using the AdamW[14] optimizer with a maximum learning rate of $3.5\times 10^{-4}$ and a weight decay of $5\times 10^{-4}$. The batch size is set to 16, and the model is trained for 180,000 iterations. A OneCycle[21] learning rate scheduler is employed to adjust the learning rate throughout training. The loss weights are set as: $\lambda_{\text{sem}}=1000,\lambda_{\text{trans}}=1,\lambda_{\text{ori}}=10$ and the iterative pose refinement is performed for $N=6$ iterations.
Metrics. We follow the evaluation metrics defined in [36, 26], including Recall@X m, Recall@X°, Absolute Position Error (APE), and Absolute Orientation Error (AOE), which jointly assess translational and rotational localization accuracy.
@@ 4.3 Comparisons with Existing Methods
We compare our proposed network HOLO with several existing methods: OrienterNet [18], U-BEV [3], MapLocNet [26], SegLocNet [36]. In addition, to further validate the effectiveness of our homography-guided strategy, we implement a pose estimation network based on cross-attention feature fusion[23, 16], namely HOLO-CA, as shown in Fig. 4. This network is also included in our comparison.